# RmGPT: Rotating Machinery Generative Pretrained Model

### v1

@@ Abstract
In industry, the reliability of rotating machinery is critical for production efficiency and safety. Current methods of Prognostics and Health Management (PHM) often rely on task-specific models, which face significant challenges in handling diverse datasets with varying signal characteristics, fault modes and operating conditions. Inspired by advancements in generative pretrained models, we propose RmGPT, a unified model for diagnosis and prognosis tasks. RmGPT introduces a novel token-based framework, incorporating Signal Tokens, Prompt Tokens, Time-Frequency Task Tokens and Fault Tokens to handle heterogeneous data within a unified model architecture. We leverage self-supervised learning for robust feature extraction and introduce a next signal token prediction pretraining strategy, alongside efficient prompt learning for task-specific adaptation. Extensive experiments demonstrate that RmGPT significantly outperforms state-of-the-art algorithms, achieving near-perfect accuracy in diagnosis tasks and exceptionally low errors in prognosis tasks. Notably, RmGPT excels in few-shot learning scenarios, achieving 92% accuracy in 16-class one-shot experiments, highlighting its adaptability and robustness. This work establishes RmGPT as a powerful PHM foundation model for rotating machinery, advancing the scalability and generalizability of PHM solutions.
keywords:
@@ 1 Introduction
Rotating machinery is a critical component in many industrial applications[1, 2], and ensuring its reliability is essential for production efficiency and safety[3]. Prognostics and Health Management (PHM) encompasses a range of methodologies aimed at assessing and predicting the present and future health status of equipment[4, 5]. The outcomes of PHM enable timely maintenance to prevent unexpected failures. Current PHM methods for fault diagnosis and prognosis often rely on task-specific models tailored to particular types of equipment and operational conditions and specific fault modes. They typically involve handcrafted features and machine learning algorithms designed to identify specific fault patterns from historical data[6, 7, 8].
@@
While effective in specific contexts, these task-specific models encounter significant challenges when applied to diverse datasets with varying signal characteristics, fault modes, and operating environments[9]. Several key difficulties arise in this context: 1) Variability in Signal Dynamics: Different types of diagnosis equipment exhibit significant variations in sensor quantity, installation position, monitored types of signals, and sampling frequencies. Modeling these diverse signals within a unified framework is highly challenging. 2) Diversity in Fault Mechanisms and Patterns: The wide range of equipment types and their differing designs and fault mechanisms complicate the definition of unified diagnosis and prognosis tasks. Each equipment type may exhibit unique fault behaviors that are difficult to generalize across other types. 3) Lack of a Foundation Model for Rotating Machinery: The differences in input signals and diagnostic tasks lead to the difficulty of having a foundation model capable of using a single set of parameters to effectively perform diagnosis and prognosis tasks across multiple types of equipment. Despite efforts by researchers to employ transfer learning techniques for domain adaptation and generalization in rotating machinery[10, 11, 12, 13, 14], the absence of such a foundation model limits knowledge transfer and sharing between different devices and hinders the application of AI techniques in real PHM scenarios[15].
Inspired by advancements in generative pretrained models like ChatGPT [16] and SAM [17], we propose RmGPT, a unified model for diagnosis and prognosis tasks in rotating machinery. RmGPT leverages self-supervised learning to capture generalizable features from vast amounts of unlabeled data and introduces a unified diagnosis and prognosis framework that integrates data and tasks from different diagnosis equipment into a generative token-based paradigm. By using a consistent model architecture and parameters, RmGPT can effectively handle the heterogeneity of rotating machinery data, encompassing various signal channels, fault categories, sampling frequencies and dataset sizes. This novel approach, illustrated in Fig. 1, not only simplifies the model development process but also enhances the model robustness and adaptability in real-world applications, demonstrating superior generalization across different types of equipment and operational conditions.
@@
1. Unified Diagnosis and Prognosis Paradigm: We propose a novel paradigm based on a semantic token space for diagnosis and prognosis. This involves using learnable Fault Tokens to capture fault prototypes of different equipment and Time-Frequency Task Tokens to represent the health status semantics of input signals. Unified diagnosis and prognosis are performed by comparing these health status semantics tokens with Fault Tokens.
2. Unified Efficient Model for Diagnosis and Prognosis: We introduce an efficient model that can adaptively handle different signal inputs. The model utilizes learnable Prompt Tokens to adaptively learn sensor semantic features from different equipment and employs an efficient Signal Token mapping method, with a dual-stage attention mechanism to efficiently and adaptively extract signal semantics. Additionally, we propose a Next Signal Token Prediction pretraining strategy to capture generalizable features from vast amounts of unlabeled data and an Efficient Prompt Learning method for high-efficiency task adaptation.
3. Superior Performance Across Diverse Datasets: Extensive experiments demonstrate that RmGPT significantly outperforms state-of-the-art (SOTA) algorithms in both diagnosis and prognosis tasks. The model achieves near-perfect accuracy in diagnosis tasks and exceptionally low errors in prognosis tasks. Additionally, RmGPT shows remarkable performance in few-shot learning scenarios, achieving high accuracy even with minimal training samples, thus validating its adaptability and robustness.
@@ 2 Problem Definition
2.1 Self-Supervised Learning for Foundation Models
@@ 2.1.1 Diagnosis and Prognosis
where $\hat{\mathcal{Y}}_{i}$ is the predicted output for dataset $D_{i}$ based on the learned tokens $\mathbf{h}_{i}$ and the same foundation model $F$ with the same parameters $\theta_{d}$. For the diagnosis task, ${Y}_{i}\in\mathbb{R}^{FC_{i}}$, the model outputs a classification among the fault categories $FC_{i}$. For the prognosis task, ${Y}_{i}\in\mathbb{R}$, the model outputs a Remaining Useful Life (RUL) prediction.
2.1.2 Prompt Learning& Finetuning
@@
2.1.3 Few-shot Learning & One-Shot Learning
@@
3 Rotating Machinery Generative Pretrained Model
In this chapter, we introduce RmGPT, as depicted in Fig. 2. RmGPT serves as a foundation model for unified diagnosis and prognosis of rotating machinery. The model employs a generative framework using tokens to represent diagnosis and prognosis tasks. It features an optimized architecture with a patch-based tokenizer and a time-channel attention transformer, ensuring efficient processing and adaptability. The training process includes a pretrained strategy focused on predicting signal tokens and a prompt-based technique for fine-tuning the model to specific tasks.
3.1 Unified Diagnosis and Prognosis Framework
To develop a foundation model that unifies the modeling of $\mathcal{X}$ and the diagnosis and prognosis $\mathcal{Y}$, we draw inspiration from large language models and multimodal frameworks, proposing the use of tokens to represent diverse data and task information. Unlike conventional approaches that require separate models for different datasets or tasks, RmGPT employs a unified architecture and set of parameters to model signals $\mathcal{X}_{i}$ from various equipment, sensor groups, frequencies, and operational conditions, along with their corresponding diagnosis and prognosis outputs $\mathcal{Y}_{i}$ for different fault categories, all within a cohesive token space.
@@ 3.1.2 Signal Tokens
3.1.3 Time-Frequency Task Tokens & Health Token
@@ 3.1.4 Fault Tokens
To implement the diagnosis and prognosis tasks, a comparison-based approach is utilized, wherein the health representations $T_{h}$ are compared against the fault tokens $T_{f}$. For diagnosis tasks, the model calculates the similarity between $T_{h}$ and each fault prototype representation in $FC_{i}$, selecting the fault category with the highest similarity as the diagnosis result. In the context of RUL prediction, the Euclidean distance between $T_{h}$ and each fault prototype representation in $FC_{i}$ is computed, with the minimum distance being transformed into the predicted RUL. This comparison-based method is driven by its flexibility and adaptability to various input scenarios, enabling the model to seamlessly handle diverse diagnostic and prognostic tasks. Furthermore, this approach provides more interpretable results, as the diagnosis and prognosis are directly linked to the specific characteristics of the identified fault mode.
3.2 Model Architecture
@@
3.2.1 Patch-Based Tokenizer
The Patch-Based Tokenizer plays a crucial role in segmenting input signals into manageable patches[18], transforming local information into semantic vectors suitable for processing by the Transformer network, as depicted in Fig. 3a. For a multivariate time series $x^{i}\in\mathbb{R}^{L_{i}\times M_{i}}$, this process involves dividing the signal into patches of length $P$ with a stride $S$, resulting in $N_{i}$ patches:
$N_{i}=\left\lfloor\frac{L_{i}-P}{S}\right\rfloor+1$ (6)
Each patch $x_{p}^{(i)}$ is then linearly transformed into a token $x_{t}^{(i)}\in\mathbb{R}^{N_{i}\times d\times M_{i}}$:
$x_{t}^{(i)}=\operatorname{Linear}(x_{p}^{(i)})$ (7)
Position encoding $W_{pos}$ is added to these tokens to incorporate temporal information, resulting in the signal tokens $T_{s}^{(i)}$:
@@ 3.2.1 Patch-Based Tokenizer
This tokenization approach significantly reduces the computational and memory requirements, which typically scale with the square of the sequence length $N$ in a Transformer network [19]. By reducing $N$ from the full length of the input sequence $L$ to approximately $\frac{L}{S}$, this method achieves a quadratic reduction in complexity. For example, in our dataset, an input sequence of length $L=2048$ is divided into patches of length $P=256$ with a stride of $S=256$, thereby significantly lowering the computational burden while preserving the semantic integrity of each token.
3.2.2 Channel-Time Attention Transformer

### v2

# RmGPT: A Foundation Model with Generative Pre-trained Transformer for Fault Diagnosis and Prognosis in Rotating Machinery

@@ Abstract
In industry, the reliability of rotating machinery is critical for production efficiency and safety. Current methods of Prognostics and Health Management (PHM) often rely on task-specific models, which face significant challenges in handling diverse datasets with varying signal characteristics, fault modes and operating conditions. Inspired by advancements in generative pretrained models, we propose RmGPT, a unified model for diagnosis and prognosis tasks. RmGPT introduces a novel generative token-based framework, incorporating Signal Tokens, Prompt Tokens, Time-Frequency Task Tokens and Fault Tokens to handle heterogeneous data within a unified model architecture. We leverage self-supervised learning for robust feature extraction and introduce a next signal token prediction pretraining strategy, alongside efficient prompt learning for task-specific adaptation. Extensive experiments demonstrate that RmGPT significantly outperforms state-of-the-art algorithms, achieving near-perfect accuracy in diagnosis tasks and exceptionally low errors in prognosis tasks. Notably, RmGPT excels in few-shot learning scenarios, achieving 82% accuracy in 16-class one-shot experiments, highlighting its adaptability and robustness. This work establishes RmGPT as a powerful PHM foundation model for rotating machinery, advancing the scalability and generalizability of PHM solutions. Code is available at: https://github.com/Pandalin98/RmGPT.
Index Terms:
@@ I Introduction
Rotating machinery is a critical component in many industrial applications[fink_potential_2020, zio_prognostics_2022, Sketching], and ensuring its reliability is essential for production efficiency and safety[liu_artificial_2018, Satellite, SAH-NET, wang2025data]. Prognostics and Health Management (PHM) encompasses a range of methodologies aimed at assessing and predicting the present and future health status of equipment[zhang2023realistic, wang2024physics, Matrix]. The outcomes of PHM enable timely maintenance to prevent unexpected failures. Current PHM methods for fault diagnosis and prognosis often rely on task-specific models tailored to particular types of equipment and operational conditions and specific fault modes. They typically involve handcrafted features and machine learning algorithms designed to identify specific fault patterns from historical data[gao_data_2024, wang_establishment_2022, wen_gru-ae-wiener_2024, Microgrid].
@@
While effective in specific contexts, these task-specific models encounter significant challenges when applied to diverse datasets with varying signal characteristics, fault modes, and operating environments[wang2023self, BasGPT]. Several key difficulties arise in this context: 1) Variability in Signal Dynamics: Different types of diagnosis equipment exhibit significant variations in sensor quantity, installation position, monitored types of signals, and sampling frequencies. Modeling these diverse signals within a unified framework is highly challenging. 2) Diversity in Fault Mechanisms and Patterns: The wide range of equipment types and their differing designs and fault mechanisms complicate the definition of unified diagnosis and prognosis tasks. Each equipment type may exhibit unique fault behaviors that are difficult to generalize across other types. 3) Lack of a Foundation Model for Rotating Machinery: The differences in input signals and diagnostic tasks lead to the difficulty of having a foundation model capable of using a single set of parameters to effectively perform diagnosis and prognosis tasks across multiple types of equipment. Despite efforts by researchers to employ transfer learning techniques for domain adaptation and generalization in rotating machinery[10473026, 10577994, 10418185, 9951150, sun2023unsupervised, 10225508], the absence of such a foundation model limits knowledge transfer and sharing between different devices and hinders the application of AI techniques in real PHM scenarios[li2023chatgpt].
Recent advancements in generative pre-trained models, such as ChatGPT [zhou2023chatgpt] and SAM [kirillov2023segment], have demonstrated the power of self-supervised learning in capturing generalizable patterns from vast amounts of unlabeled data. Inspired by these breakthroughs, we propose RmGPT, a generative foundation model designed to adapt to diverse fault diagnosis and prognosis tasks in rotating machinery. Unlike traditional discriminative models that rely on manually crafted features or task-specific finetuning, RmGPT takes a generative approach, modeling the machinery’s health status by learning to predict and generate the signal patterns from raw data.
RmGPT introduces a unified token-based framework that incorporates Signal Tokens, Prompt Tokens, Time-Frequency Task Tokens, and Fault Tokens. This design enables the model to autonomously adapt to varying signal characteristics and fault modes across different equipment types. By learning to generate relevant health information through these tokens, RmGPT produces a robust representation of machinery health. Although inspired by generative transformer architectures, RmGPT is not a language model. It is specifically designed for multivariate time-series signals and operates entirely on continuous sensor data without using textual inputs or natural language supervision. Leveraging self-supervised learning and next-token prediction, the model efficiently captures the underlying structures in diverse signal data, ensuring high generalization across datasets with minimal labeled data. As shown in Fig. 1, this approach empowers RmGPT to generate consistent health representations that are seamlessly integrated into downstream tasks, making it highly adaptable and capable of handling different equipment and operational conditions.
@@
1. Unified Diagnosis and Prognosis Paradigm: We propose a novel generative paradigm based on a semantic token space for diagnosis and prognosis tasks. This paradigm leverages learnable Fault Tokens to capture fault prototypes from different equipment types and Time-Frequency Task Tokens to represent the health status semantics of input signals. By comparing these generative health status semantics tokens with Fault Tokens, RmGPT performs unified diagnosis and prognosis across varying equipment and fault conditions.
2. Unified Efficient Model for Adaptive Signal Processing: We introduce an efficient generative model that can adaptively handle varying signal inputs from different equipment types. The model utilizes learnable Prompt Tokens to capture sensor-specific semantics and employs an efficient Signal Token mapping method. A dual-stage attention mechanism is used to generatively extract and refine signal features. Additionally, we propose a Next Signal Token Prediction pretraining strategy, which enables RmGPT to capture generalizable features from vast amounts of unlabeled data, and an Efficient Prompt Learning method for rapid adaptation to new tasks.
3. Unified Superior Performance Across Diverse Datasets: Extensive experiments demonstrate that RmGPT outperforms state-of-the-art (SOTA) algorithms in both diagnosis and prognosis tasks. The model achieves near-perfect accuracy in diagnosis tasks and exceptionally low errors in prognosis tasks. Moreover, RmGPT excels in few-shot learning scenarios, maintaining high accuracy even with minimal training samples, thus validating its generative robustness and adaptability across diverse datasets and equipment types.
@@ II Problem Definition
II-A Definition and Characteristics of Foundation Models in PHM
Foundation models have recently emerged as a dominant paradigm in natural language processing and computer vision, where a single, pre-trained model can be adapted to various downstream tasks with minimal task-specific adjustment. However, applying this concept directly to Prognostics and Health Management (PHM)—particularly in the context of rotating machinery—requires careful reinterpretation, due to the structured, heterogeneous, and domain-specific nature of industrial data.
In this work, we define a PHM foundation model as:
A general-purpose model $F$ with shared parameters $\theta$, trained via self-supervised learning on heterogeneous time-series datasets $\{D_{i}\}$, capable of producing task-relevant representations $\mathbf{h}_{i}=F(\mathcal{X}_{i};\theta)$ and supporting diverse diagnostic and prognostic tasks across different equipment and conditions without modifying the model architecture.
Compared to traditional approaches where each dataset $D_{i}=(\mathcal{X}_{i},\mathcal{Y}_{i})$ is handled by an independent model $F_{i}$ with task-specific parameters $\theta_{i}$, a PHM foundation model aims to unify representation learning and task modeling through a consistent parameter set $\theta$ shared across all datasets.
We summarize the key properties of a PHM foundation model as follows:
  • Cross-domain representation learning: The model is trained on a diverse collection of datasets $\{D_{i}\}$ with varying input dimensions, sampling frequencies, fault types, and operating conditions, and is expected to learn generalizable representations $\mathbf{h}_{i}$ from unlabeled data.
  • Multi-task flexibility: The learned representations should support a range of PHM tasks (e.g., Fault Diagnosis, Remaining Useful Life (RUL) prediction) in a consistent framework without requiring redesign of task-specific modules.
  • Architecture-level unification: The model maintains a consistent network architecture across datasets and tasks, facilitating adaptation to new application scenarios without structural modification.
This reformulation of the foundation model paradigm in PHM provides a theoretical basis for constructing general-purpose models that unify diagnosis and prognosis across rotating machinery datasets, while offering scalability to broader industrial applications.
II-B Self-Supervised Learning for Foundation Models
@@ II-B1 Diagnosis and Prognosis
where $\hat{\mathcal{Y}}_{i}$ is the predicted output for dataset $D_{i}$ based on the learned tokens $\mathbf{h}_{i}$ and the same foundation model $F$ with the same parameters $\theta_{d}$. For the diagnosis task, ${Y}_{i}\in\mathbb{R}^{FC_{i}}$, the model outputs a classification among the fault categories $FC_{i}$. For the prognosis task, ${Y}_{i}\in\mathbb{R}$, the model outputs a RUL prediction.
II-B2 Prompt Learning& Finetuning
@@
II-B3 Few-shot Learning & One-Shot Learning
@@
III Rotating Machinery Generative Pretrained Transformer
In this chapter, we introduce RmGPT, as depicted in Fig. 2. RmGPT serves as a foundation model for unified diagnosis and prognosis of rotating machinery. Although RmGPT draws structural inspiration from large generative models, it is not a language model. Instead, it is specifically designed to process multivariate time-series signals in industrial scenarios and does not rely on textual inputs or language modeling objectives. The model employs a generative framework using tokens to represent diagnosis and prognosis tasks. It features an optimized architecture with a patch-based tokenizer and a channel-time attention transformer, ensuring efficient processing and adaptability. The training process includes a pretrained strategy focused on predicting signal tokens and a prompt-based technique for finetuning the model to specific tasks.
III-A Unified Diagnosis and Prognosis Framework
To develop a foundation model that unifies the modeling of $\mathcal{X}$ and the diagnosis and prognosis $\mathcal{Y}$, we draw inspiration from large language models and multimodal frameworks, proposing the use of tokens to represent diverse data and task information. Unlike conventional approaches that require separate models for different datasets or tasks, RmGPT employs a unified architecture with a consistent set of parameters to model signals $\mathcal{X}_{i}$ from various equipment, sensor groups, frequencies, and operational conditions, alongside their corresponding diagnosis and prognosis outputs $\mathcal{Y}_{i}$ for different fault categories. The model generates outputs based on health status semantic tokens, all within a cohesive token space.
@@ III-A2 Signal Tokens
III-A3 Time-Frequency Task Tokens & Health Token
@@ III-A4 Fault Tokens
To enhance the interpretability of the model’s internal inference mechanism, as illustrated in Fig. 2(c), we provide a detailed explanation of how different types of tokens—particularly the Time-Frequency Task Tokens ($T_{t}$) and Fault Tokens ($T_{f}$)—interact during the inference process to support fault diagnosis and prognosis decisions.
Prior to the channel-time attention layers, we concatenate the Prompt Tokens ($T_{p}\in\mathbb{R}^{l_{p}\times M\times d}$), Signal Tokens ($T_{s}\in\mathbb{R}^{l_{s}\times M\times d}$), and Time-Frequency Task Tokens ($T_{t}\in\mathbb{R}^{l_{t}\times M\times d}$) along the sequence dimension to form the model’s input:
$T_{\text{in}}=T_{p}\ \|\ T_{s}\ \|\ T_{t}\in\mathbb{R}^{N\times M\times d},$
where $N=l_{p}+l_{s}+l_{t}$ is the total sequence length. This unified token sequence is fed into the backbone encoder for joint modeling of channel and temporal dependencies:
$T_{\text{out}}=F_{\text{enc}}(T_{\text{in}})\in\mathbb{R}^{N\times M\times d}.$
We then extract the output embeddings corresponding to the Time-Frequency Task Tokens from the last $l_{t}$ positions of $T_{\text{out}}$, and compute the Health Token $T_{h}\in\mathbb{R}^{M\times d}$ by applying mean pooling over the sequence dimension:
$T_{h}=\frac{1}{l_{t}}\sum_{i=N-l_{t}+1}^{N}T_{\text{out}}^{(i)},$
where $T_{\text{out}}^{(i)}\in\mathbb{R}^{M\times d}$ denotes the $i$-th token in the output sequence. This compact representation $T_{h}$ captures the aggregated temporal and frequency-domain health semantics of the system, and is subsequently compared with learned fault prototypes for diagnosis and prognosis.
Following the channel-time attention layers, the input signal is first converted into a series of Signal Tokens ($T_{s}\in\mathbb{R}^{l_{s}\times M\times d}$), where $l_{s}$ is the token length, $M$ is the number of sensor channels, and $d$ is the embedding dimension. These tokens are further processed and transformed into Time-Frequency Task Tokens ($T_{t}\in\mathbb{R}^{l_{t}\times M\times d}$), which capture both time-domain and frequency-domain characteristics of the input signals. The $T_{t}$ tokens are aggregated through attention mechanisms into Health Tokens ($T_{h}\in\mathbb{R}^{M\times d}$), serving as compact latent representations of the system’s current health state.
During inference, $T_{h}$ is compared against a set of learnable Fault Tokens $T_{f}=\{T_{f}^{(1)},T_{f}^{(2)},\dots,T_{f}^{(C)}\}$, where each $T_{f}^{(c)}\in\mathbb{R}^{M\times d}$ represents the prototype embedding of fault class $c$, and $C$ is the total number of fault categories. For classification tasks (i.e., fault diagnosis), the model computes the similarity between $T_{h}$ and each $T_{f}^{(c)}$ using a vector similarity function $\text{Sim}(\cdot,\cdot)$ (e.g., cosine similarity). The predicted class label $\hat{y}_{\text{diag}}$ is assigned to the class with the highest similarity:
$\hat{y}_{\text{diag}}=\arg\max_{c}\ \text{Sim}(T_{h},T_{f}^{(c)})$ (6)
For regression tasks (i.e., Remaining Useful Life prediction), the model measures the Euclidean distance between $T_{h}$ and each $T_{f}^{(c)}$, and maps the smallest distance to an estimated RUL value $\hat{y}_{\text{rul}}$ via a transformation function $g(\cdot)$:
$\hat{y}_{\text{rul}}=g\left(\min_{c}\|T_{h}-T_{f}^{(c)}\|\right)$ (7)
Here, $g(\cdot)$ is a task-specific regression function learned during training. This inference strategy enables the unified token representation to support both classification and regression tasks in a seamless and consistent manner.
By comparing the semantic health state (captured by $T_{t}$ and aggregated into $T_{h}$) with class-level prototypes ($T_{f}$), the model achieves generalization across multiple fault types and signal domains. This prototype-driven matching mechanism, as depicted in Fig. 2(c), reflects the core design philosophy of RmGPT: a unified, interpretable, and extensible inference framework built upon token-based semantic alignment in a latent space.
III-B Model Architecture
@@
III-B1 Patch-Based Tokenizer
The Patch-Based Tokenizer plays a crucial role in segmenting input signals into manageable patches [wang_incorporating_2024], transforming local information into semantic vectors suitable for processing by the Transformer network, as illustrated in Fig. 3a. Given a multivariate time series $x^{(i)}\in\mathbb{R}^{L_{i}\times M_{i}}$, where $L_{i}$ is the sequence length and $M_{i}$ is the number of sensor channels, we first segment each univariate channel independently into $l_{s}^{(i)}$ temporal patches of fixed length $P$ and stride $S$, where:
$l_{s}^{(i)}=\left\lfloor\frac{L_{i}-P}{S}\right\rfloor+1$ (8)
This converts each 1D input sequence of length $L_{i}$ into a sequence of $l_{s}^{(i)}$ localized subsequences (patches), preserving the temporal continuity within each segment.
Each patch is then projected into a $d$-dimensional latent space using a shared linear transformation across channels, resulting in a compact embedding. By applying this process to all channels and stacking the results, we obtain a token tensor of shape:
$x_{t}^{(i)}=\operatorname{Linear}(x_{p}^{(i)})\in\mathbb{R}^{l_{s}^{(i)}\times M_{i}\times d}$ (9)
where the first dimension indexes temporal patches, the second dimension corresponds to sensor channels, and the third dimension is the latent embedding.
To encode temporal order, a positional encoding matrix $W_{pos}\in\mathbb{R}^{l_{s}^{(i)}\times d}$ is added to each channel’s token sequence via broadcasting:
@@ III-B1 Patch-Based Tokenizer
The resulting matrix $T_{s}^{(i)}\in\mathbb{R}^{l_{s}^{(i)}\times M_{i}\times d}$ serves as the input to the Transformer encoder.
This tokenization strategy offers two key advantages. First, it significantly reduces the computational and memory cost of self-attention, which scales quadratically with sequence length [nie2022time]. By converting the input from $L_{i}$ time points to approximately $L_{i}/S$ patch tokens, the model achieves substantial efficiency gains—for example, a sequence of length $L=2048$ with $P=S=256$ results in only $l_{s}=8$ tokens per channel. Second, representing each patch as a semantically meaningful unit allows the model to capture richer local temporal dynamics and structural patterns, leading to more robust and generalizable representations.
III-B2 Channel-Time Attention Transformer